Example 3-1

During 20--, Rachael Parkins, president of Mathieson Company, was paid a semimonthly salary of $7,600. Compute the amount of FICA taxes that should be withheld from her. If an amount is zero, enter "0".

OASDI HI
a. 9th paycheck $ $
b. 18th paycheck $ $
c. 24th paycheck $ $
d. If Rachael Parkin's year-to-date earnings as of her 24th paycheck
     are $182,400, and her year-end bonus is $100,000, how much is
     the additional 0.9% HI tax withheld?
Answer #1
OASDI HI
a. 9th paycheck 471.20 110.20
b. 18th paycheck 229.40 110.20
c. 24th paycheck 0 110.20
d. Additional HI tax withheld 741.60
Workings:
The FICA OASDI tax limit is $132900
OASDI tax:
9th paycheck OASDI 471.20 =7600*6.2%
Cumulative earnings till 17th paycheck 129200 =7600*17
18th paycheck OASDI 229.40 =(132900-129200)*6.2%
HI tax 110.20 =7600*1.45%
Additional HI tax is payable for wages above $200000
Additional HI tax withheld 741.60 =(182400+100000-200000)*0.9%